-2 { -9(-12)+ 2 }
3 { -4(-16) }
1
Expert's answer
2014-05-28T13:16:52-0400
-2 { -9(-12)+ 2 }=-2{108+2}=-2*110=-220;
3 { -4(-16) }=3*{64}=3*64=192